OpenStudy (anonymous):

how do you write in standard form, an equation of the line passing through (-2,-2) with a slope of -8, how do i find the y-intersect?

hero (hero):

Hint: y = mx + b (x,y) = (-2,-2) m = -8 b = y-intercept

OpenStudy (anonymous):

You can use the point slope form formula to find the y intersect. Use y =mx+b To find b, which is our y-intersect substitute our point (-2,-2) into our equation y=mx+b -2=-8(-2) + b -2=16 +b -2-16=b -18=b b = -18 so our y-intersect that passes through the point (-2,-2) with a slope of -8 would be (0,-18) Now we have b we can turn it into standard form y = -8x + -18 8x + y = -18

OpenStudy (anonymous):

thank you